# Feature learning in finite-width Bayesian deep linear networks with multiple outputs and convolutional layers

### v1

@@ III.2 Posterior distribution and predictor’s statistics in the case of squared error loss function
$p_{\mathrm{pred}}(\mathbf{s}_{0}|\mathbf{y}_{1:P},\mathbf{x}_{0},X)=\int_{(\mathcal{S}^{+}_{D})^{L}}\frac{e^{-\frac{1}{2}(\mathbf{s}_{0}-\mathbf{m}_{0})^{\top}(\Sigma_{00}-\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{DP})^{-1}\Sigma_{01}^{\top})^{-1}(\mathbf{s}_{0}-\mathbf{m}_{0})}}{(2\pi)^{\frac{D}{2}}\det(\Sigma_{00}-\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{DP})^{-1}\Sigma_{01}^{\top})^{\frac{1}{2}}}\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L}|\mathbf{y}_{1:P},\beta)$
@@
$\Sigma_{00}=Q^{(L)}\otimes\frac{\mathbf{x}_{0}^{\top}\mathbf{x}_{0}}{N_{0}\lambda^{*}},\quad\Sigma_{01}=Q^{(L)}\otimes\frac{\mathbf{x}_{0}^{\top}X}{N_{0}\lambda^{*}},\quad\Sigma_{11}=Q^{(L)}\otimes\frac{X^{\top}X}{N_{0}\lambda^{*}}\quad\mathbf{m}_{0}=\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{DP})^{-1}\\ \mathbf{y}_{1:P},$
$\qquad\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L}|\mathbf{y}_{1:P},\beta):=\frac{e^{-\frac{1}{2}\Phi_{\beta}(Q_{1}\dots Q_{L},\mathbf{y}_{1:P})}\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L})}{\int_{(\mathcal{S}^{+}_{D})^{L}}e^{-\frac{1}{2}\Phi_{\beta}(Q_{1}\dots Q_{L},\mathbf{y}_{1:P})}\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L})}$
@@
$p_{\mathrm{pred}}(s_{0}|\mathbf{y}_{1:P},\mathbf{x}_{0},X)=\int_{(\mathcal{S}^{+}_{N_{0}})^{L}}\frac{e^{-\frac{1}{2}(\Sigma_{00}-\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{P})^{-1}\Sigma_{01}^{\top})^{-1}(s_{0}-m_{0})^{2}}}{(2\pi)^{\frac{1}{2}}\det(\Sigma_{00}-\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{P})^{-1}\Sigma_{01}^{\top})^{\frac{1}{2}}}\mathcal{Q}_{L,C}(dQ_{1}\dots dQ_{L}|\mathbf{y}_{1:P},\beta)$
@@
$\begin{split}&\Sigma_{00}:=\sum_{r,s}\mathcal{T}(Q_{1},\dots,Q_{1})_{rs}\sum_{a_{0}}\frac{x^{0}_{a_{0},r}x^{0}_{a_{0},s}}{\lambda^{*}C_{0}},\quad\Sigma_{01}:=\Big{[}\sum_{r,s}\mathcal{T}(Q_{1},\dots,Q_{1})_{rs}\sum_{a_{0}}\frac{x^{\mu}_{a_{0},r}x^{0}_{a_{0},s}}{\lambda^{*}C_{0}}\Big{]}_{\mu=1,\dots,P}\\ &\Sigma_{11}:=\Big{[}\sum_{r,s}\mathcal{T}(Q_{1},\dots,Q_{1})_{rs}\sum_{a_{0}}\frac{x^{\mu}_{a_{0},r}x^{\nu}_{a_{0},s}}{\lambda^{*}C_{0}}\Big{]}_{\mu,\nu=1,\dots,P},\quad m_{0}:=\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{P})^{-1}\mathbf{y}_{1:P},\\ &\mathcal{Q}_{L,C}(dQ_{1}\dots dQ_{L}|\mathbf{y}_{1:P},\beta):=\frac{e^{-\frac{1}{2}\Phi_{\beta}(Q_{1}\dots Q_{L},\mathbf{y}_{1:P})}\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L})}{\int_{(\mathcal{S}^{+}_{N_{0}})^{L}}e^{-\frac{1}{2}\Phi_{\beta}(Q_{1}\dots Q_{L},\mathbf{y}_{1:P})}\mathcal{Q}_{L,C}(dQ_{1}\dots dQ_{L})}\\ \end{split}$
@@
III.3 Quantitative description of the feature learning infinite-width regime
@@
Following the literature [3, 6, 30, 31, 10], we recall that the easiest way to escape the lazy-training infinite-width limit is to consider the so-called mean-field parametrization, which corresponds to the following rescaling of the loss and output functions:
@@
Roughly speaking, if $V_{N}\in\mathbb{R}^{D}$ satisfies an LDP with rate function $I$ then $P\{V_{N}\in dx\}=P_{N}(dx)\simeq e^{-NI(x)}dx$. The formal definition [59] is as follows: a sequence of random elements $(V_{N})_{N}$ taking values in a metric space $(\mathbb{V},d_{V})$ (or equivalently the sequence of their laws, say $P_{N}$) satisfies an LDP with rate function $I$ if
[…]
and
[…]
@@
As a corollary, using the well-known contraction principle (see, e.g., Theorem 4.2.1 in [59] ), if $(Q_{1},\dots,Q_{L})$ has law $\mathcal{Q}_{L,N}$, then $Q^{(L)}(Q_{1},\dots,Q_{L})$ satisfies an LDP on $\mathcal{S}^{+}_{D}$ with rate function
@@
The rate function now explicitly depends on the training inputs and labels, showing that the measure over the Wishart ensembles this time concentrates around non-trivial data-dependent solutions, which quantitatively characterize feature learning in this limit (see also Discussion in Section IV for a comparison with the proportional limit studied in [19, 36, 34]). We observe that the result in Proposition 6 is complementary to the recent investigation in Ref. [60], where the feature learning infinite-width limit of DLNs under gradient descent dynamics was characterized.
@@
Propositions 5 and 6 hold for the C-DLN with $\mathcal{Q}_{L,N}$ and $\mathcal{Q}^{\circ}_{L,N}$ replaced by $\mathcal{Q}_{L,C}$ and $\mathcal{Q}^{\circ}_{L,C}$. In this case $C\to+\infty$ and $\Phi_{\beta}^{\circ}$ has the same form given in (17) but $\Sigma_{11}$ is defined as in Proposition 4.
IV Discussion: from Kernel renormalization to rigorous Bayesian statistics
In concluding, we discuss how our derivations provide a way to interpret the recent statistical-physics literature on Bayesian deep learning theory in the proportional limit ($P,N_{\ell}\to\infty$, fixed $\alpha_{\ell}=P/N_{\ell}$) [19, 32, 36, 54].
Ref. [34] provides the first notable results that partly address this issue. However, the representation of the Bayesian model evidence (partition function) in terms of Meijer G-functions hides the possibility of a direct comparison with the previously-established physics results. On the contrary, our explicit non-asymptotic integral representation of the prior delivers a precise one-to-one mapping between the idea of kernel renormalization introduced by physicists [19, 36, 54] and the Wishart ensembles that appear in our rigorous derivation.
@@ IV Discussion: from Kernel renormalization to rigorous Bayesian statistics
$Z_{\mathbf{y}_{1:P},\beta}\propto\int_{0}^{\infty}\prod_{\ell=1}^{L}dQ_{\ell}\,e^{-\sum_{\ell=1}^{L}\left[\frac{N_{\ell}}{2}Q_{\ell}-\left(\frac{N_{\ell}}{2}-1\right)\log Q_{\ell}\right]-\frac{1}{2}\log\det\beta\left[\frac{\mathbb{1}}{\beta}+\mathcal{K}_{\mathrm{FC}}(Q_{1},\dots,Q_{L})\right]-\frac{1}{2}\mathbf{y}_{1:P}^{\top}\left[\frac{\mathbb{1}}{\beta}+\mathcal{K}_{\mathrm{FC}}(Q_{1},\dots,Q_{L})\right]^{-1}\mathbf{y}_{1:P}}\,,$ (18)
@@
which is exact up to an overall unimportant constant. Note that here the first two terms in the exponential arise from the probability measure over the scalar variables $Q_{\ell}$, while the third and fourth data-dependent terms are a consequence of the Gaussian integration over the output variables. This partition function can be compared with the result for non-linear networks found in the proportional limit in [36]. It turns out that if we restrict our focus to linear NNGP kernels in [36], we find an effective action equivalent to the exponent in Eq. (18) at leading order in $N_{\ell}$. Moreover, it is reasonable to expect that the third and fourth terms in Eq. (18) scale linearly with $P$, under proper hypotheses on the training data (note that this requirement is similar to the one in [34]). This observation suggests that it is legitimate to employ large deviations techniques (or the more powerful saddle-point method) in the proportional limit, and it provides the interpretation mentioned at the beginning of this section, since one can check that the saddle-point equations derived from Eq. (18) are exactly the same as those found in [19].
@@
By a line of reasoning similar to that exposed above for $D=1$, we can also interpret the (heuristic) kernel shape renormalization pointed out in [19, 36] for one-hidden-layer fully-connected architectures with multiple outputs, as well as the local kernel renormalization proposed to describe shallow CNNs in [54]. If we consider for instance the effective action derived in Ref. [54], we again recognize that the first two terms arise from the probability measure of a Wishart ensemble, whereas the last two data-dependent terms are a consequence of the integration of the output variables. Even in this case, one can check a precise equivalence with the non-asymptotic partition function that one finds in the CNN case using our result for the prior in Proposition 3 for $L=1$.
@@
Finally, we point out that the feature learning infinite-width limit investigated in Proposition 6 shares many similarities with the proportional limit studied in the physics literature. In both cases (but for different reasons), the measure over the Wishart ensembles concentrates over data-dependent solutions of a rate function (or a saddle-point effective action). Furthermore, the three terms of the rate function in Proposition 6 also appear in the effective action found in [36], which is known to reproduce the saddle-point equations firstly identified in [19] for the linear case.
Acknowledgements
@@
P.R. is supported by ##\##NEXTGENERATIONEU (NGEU) and funded by the Ministry of University and Research (MUR), National Recovery and Resilience Plan (NRRP), project MNESYS (PE0000006) “A Multiscale integrated approach to the study of the nervous system in health and disease” (DN. 1553 11.10.2022). F.B. is partially supported by the MUR - PRIN project “Discrete random structures for Bayesian learning and prediction” no. 2022CLTYP4.
@@ A.1 Preliminary facts
We collect here some properties of multivariate normals, matrix normals and Wishart distributions to be used later. See [61] for details. In what follows $\operatorname{vec}(A)$ denotes the stacking of the columns of a matrix $A$ to form a vector.
@@
    Equivalence with the multivariate normal. $Z\sim\mathcal{MN}(0,\Sigma_{1},\Sigma_{2})$ if and only if $\operatorname{vec}(Z)\sim\mathcal{N}(0,\Sigma_{1}\otimes\Sigma_{2})$.
@@
    Versions of this formula can be found in literature as Ingham-Siegel integrals [62, 63] (see [64] for an extension to Hermitian matrices); in this generality, the statement can be derived by Proposition 8.3 and its proof in [55]. In particular the formula above holds for $\alpha>0$ and $S=\mathbb{1}_{D}$ and $C$ symmetric and $\geq 0$, showing that when $\mathbf{Z}_{j}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,C)$, then

### v2

@@ 3.2 Posterior distribution and predictor’s statistics in the case of squared error loss function
$\begin{split}&p_{\mathrm{pred}}(\mathbf{s}_{0}|\mathbf{y}_{1:P},\mathbf{x}_{0},X)\\ &\quad=\int_{(\mathcal{S}^{+}_{D})^{L}}\!\!\frac{e^{-\frac{1}{2}(\mathbf{s}_{0}-\mathbf{m}_{0})^{\top}(\Sigma_{00}-\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{DP})^{-1}\Sigma_{01}^{\top})^{-1}(\mathbf{s}_{0}-\mathbf{m}_{0})}}{(2\pi)^{\frac{D}{2}}\det(\Sigma_{00}-\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{DP})^{-1}\Sigma_{01}^{\top})^{\frac{1}{2}}}\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L}|\mathbf{y}_{1:P},\beta)\\ \end{split}$
@@
$\begin{split}&\Sigma_{00}=\frac{\mathbf{x}_{0}^{\top}\mathbf{x}_{0}}{N_{0}\lambda^{*}}\otimes Q^{(L)},\quad\Sigma_{01}=\frac{\mathbf{x}_{0}^{\top}X}{N_{0}\lambda^{*}}\otimes Q^{(L)},\quad\Sigma_{11}=\frac{X^{\top}X}{N_{0}\lambda^{*}}\otimes Q^{(L)}\\ &\mathbf{m}_{0}=\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{DP})^{-1}\mathbf{y}_{1:P},\\ \end{split}$
$\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L}|\mathbf{y}_{1:P},\beta):=\frac{e^{-\frac{1}{2}\Phi_{\beta}(Q_{1}\dots Q_{L},\mathbf{y}_{1:P})}\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L})}{\int_{(\mathcal{S}^{+}_{D})^{L}}e^{-\frac{1}{2}\Phi_{\beta}(Q_{1}\dots Q_{L},\mathbf{y}_{1:P})}\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L})}$
@@
$\begin{split}&p_{\mathrm{pred}}(s_{0}|\mathbf{y}_{1:P},\mathbf{x}_{0},X)\\ &\quad=\int_{(\mathcal{S}^{+}_{N_{0}})^{L}}\frac{e^{-\frac{1}{2}(\Sigma_{00}-\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{P})^{-1}\Sigma_{01}^{\top})^{-1}(s_{0}-m_{0})^{2}}}{(2\pi)^{\frac{1}{2}}\det(\Sigma_{00}-\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{P})^{-1}\Sigma_{01}^{\top})^{\frac{1}{2}}}\mathcal{Q}_{L,C}(dQ_{1}\dots dQ_{L}|\mathbf{y}_{1:P},\beta)\\ \end{split}$
@@
$\begin{split}&\Sigma_{00}:=\sum_{r,s}\mathcal{T}(Q_{1},\dots,Q_{1})_{rs}\sum_{a_{0}}\frac{x^{0}_{a_{0},r}x^{0}_{a_{0},s}}{\lambda^{*}C_{0}},\\ &\Sigma_{01}:=\Big{[}\sum_{r,s}\mathcal{T}(Q_{1},\dots,Q_{1})_{rs}\sum_{a_{0}}\frac{x^{\mu}_{a_{0},r}x^{0}_{a_{0},s}}{\lambda^{*}C_{0}}\Big{]}_{\mu=1,\dots,P}\\ &\Sigma_{11}:=\Big{[}\sum_{r,s}\mathcal{T}(Q_{1},\dots,Q_{1})_{rs}\sum_{a_{0}}\frac{x^{\mu}_{a_{0},r}x^{\nu}_{a_{0},s}}{\lambda^{*}C_{0}}\Big{]}_{\mu,\nu=1,\dots,P},\\ &m_{0}:=\Sigma_{01}(\Sigma_{11}+\beta^{-1}\mathbb{1}_{P})^{-1}\mathbf{y}_{1:P},\\ &\mathcal{Q}_{L,C}(dQ_{1}\dots dQ_{L}|\mathbf{y}_{1:P},\beta):=\frac{e^{-\frac{1}{2}\Phi_{\beta}(Q_{1}\dots Q_{L},\mathbf{y}_{1:P})}\mathcal{Q}_{L,N}(dQ_{1}\dots dQ_{L})}{\int_{(\mathcal{S}^{+}_{N_{0}})^{L}}e^{-\frac{1}{2}\Phi_{\beta}(Q_{1}\dots Q_{L},\mathbf{y}_{1:P})}\mathcal{Q}_{L,C}(dQ_{1}\dots dQ_{L})}\\ \end{split}$
@@
3.3 Quantitative description of the feature learning infinite-width regime
@@
Following the literature (Chizat et al., 2019; Mei et al., 2018; Geiger et al., 2021, 2020; Yang and Hu, 2021), we recall that the easiest way to escape the lazy-training infinite-width limit is to consider the so-called mean-field parametrization, which corresponds to the following rescaling of the loss and output functions:
@@
Roughly speaking, if $V_{N}\in\mathbb{R}^{D}$ satisfies an LDP with rate function $I$ then $P\{V_{N}\in dx\}=P_{N}(dx)\simeq e^{-NI(x)}dx$. The formal definition (see, e.g., Dembo and Zeitouni, 2010) is as follows: a sequence of random elements $(V_{N})_{N}$ taking values in a metric space $(\mathbb{V},d_{V})$ (or equivalently the sequence of their laws, say $P_{N}$) satisfies an LDP with rate function $I$ if
$\limsup_{N\to+\infty}\frac{1}{N}\log\Big{(}\mathbb{P}\{V_{N}\in C\}\Big{)}=\limsup_{N\to+\infty}\frac{1}{N}\log\Big{(}P_{N}(C)\Big{)}\leq-\inf_{v\in C}I(v)$
for any closed set $C\subset\mathbb{V}$ and
$\liminf_{N\to+\infty}\frac{1}{N}\log\Big{(}\mathbb{P}\{V_{N}\in O\}\Big{)}=\liminf_{N\to+\infty}\frac{1}{N}\log\Big{(}P_{N}(O)\Big{)}\geq-\inf_{v\in O}I(v)$
for any open set $O\subset\mathbb{V}$.
@@
As a corollary, using the well-known contraction principle (see, e.g., Theorem 4.2.1 in Dembo and Zeitouni, 2010), if $(Q_{1},\dots,Q_{L})$ has law $\mathcal{Q}_{L,N}$, then $Q^{(L)}(Q_{1},\dots,Q_{L})$ satisfies an LDP on $\mathcal{S}^{+}_{D}$ with rate function
@@
The rate function now explicitly depends on the training inputs and labels, showing that the measure over the Wishart ensembles this time concentrates around non-trivial data-dependent solutions, which quantitatively characterize feature learning in this limit (see also Discussion in Section 4 for a comparison with the proportional limit studied in Li and Sompolinsky, 2021; Pacelli et al., 2023; Hanin and Zlokapa, 2023). We observe that the result in Proposition 10 is complementary to the recent investigation in Chizat et al. (2022), where the feature learning infinite-width limit of DLNs under gradient descent dynamics was characterized.
@@
Propositions 8 and  10 hold for the C-DLN with $\mathcal{Q}_{L,N}$ and $\mathcal{Q}^{\circ}_{L,N}$ replaced by $\mathcal{Q}_{L,C}$ and $\mathcal{Q}^{\circ}_{L,C}$. In this case $C\to+\infty$ and $\Phi_{\beta}^{\circ}$ has the same form given in (17) but $\Sigma_{11}$ is defined as in Proposition 7.
4 Discussion: from Kernel renormalization to rigorous Bayesian statistics
In concluding, we discuss how our derivations provide a way to interpret the recent statistical-physics literature on Bayesian deep learning theory in the proportional limit ($P,N_{\ell}\to\infty$, fixed $\alpha_{\ell}=P/N_{\ell}$) such as Li and Sompolinsky (2021, 2022); Pacelli et al. (2023); Aiudi et al. (2023).
Hanin and Zlokapa (2023) provides the first notable results that partly address this issue. However, the representation of the Bayesian model evidence (partition function) in terms of Meijer G-functions hides the possibility of a direct comparison with the previously-established physics results. On the contrary, our explicit non-asymptotic integral representation of the prior delivers a precise one-to-one mapping between the idea of kernel renormalization introduced by physicists (Li and Sompolinsky, 2021; Pacelli et al., 2023; Aiudi et al., 2023) and the Wishart ensembles that appear in our rigorous derivation.
@@ 4 Discussion: from Kernel renormalization to rigorous Bayesian statistics
$Z_{\mathbf{y}_{1:P},\beta}\propto\int_{0}^{\infty}\prod_{\ell=1}^{L}dQ_{\ell}\,e^{-\sum_{\ell=1}^{L}\left[\frac{N_{\ell}}{2}Q_{\ell}-\left(\frac{N_{\ell}}{2}-1\right)\log Q_{\ell}\right]-\frac{1}{2}\phi_{\beta}(Q_{1},\dots,Q_{L},\mathbf{y}_{1:P})}\,,$ (18)
@@
which is exact up to an overall unimportant constant. Note that here the first two terms in the exponential arise from the probability measure over the scalar variables $Q_{\ell}$, while the third and fourth data-dependent terms are a consequence of the Gaussian integration over the output variables. This partition function can be compared with the result for non-linear networks found in the proportional limit in Pacelli et al. (2023). It turns out that if we restrict our focus to linear NNGP kernels in Pacelli et al. (2023), we find an effective action equivalent to the exponent in Eq. (18) at leading order in $N_{\ell}$. Moreover, it is reasonable to expect that the third and fourth terms in Eq. (18) scale linearly with $P$, under proper hypotheses on the training data (note that this requirement is similar to the one in Hanin and Zlokapa, 2023). This observation suggests that it is legitimate to employ large deviations techniques (or the more powerful saddle-point method) in the proportional limit, and it provides the interpretation mentioned at the beginning of this section, since one can check that the saddle-point equations derived from Eq. (18) are exactly the same as those found in Li and Sompolinsky (2021).
@@
By a line of reasoning similar to that exposed above for $D=1$, we can also interpret the (heuristic) kernel shape renormalization pointed out in Li and Sompolinsky (2021) and Pacelli et al. (2023) for one-hidden-layer fully-connected architectures with multiple outputs, as well as the local kernel renormalization proposed to describe shallow CNNs in Aiudi et al. (2023). If we consider for instance the effective action derived in  Aiudi et al. (2023), we again recognize that the first two terms arise from the probability measure of a Wishart ensemble, whereas the last two data-dependent terms are a consequence of the integration of the output variables. Even in this case, one can check a precise equivalence with the non-asymptotic partition function that one finds in the CNN case using our result for the prior in Proposition 6 for $L=1$.
@@
Finally, we point out that the feature learning infinite-width limit investigated in Proposition 10 shares many similarities with the proportional limit studied in the physics literature. In both cases (but for different reasons), the measure over the Wishart ensembles concentrates over data-dependent solutions of a rate function (or a saddle-point effective action). Furthermore, the three terms of the rate function in Proposition 10 also appear in the effective action found in Pacelli et al. (2023), which is known to reproduce the saddle-point equations firstly identified in Li and Sompolinsky (2021) for the linear case.
Acknowledgments and Disclosure of Funding
@@
AP.R. is supported by ##\##NEXTGENERATIONEU (NGEU) and funded by the Ministry of University and Research (MUR), National Recovery and Resilience Plan (NRRP), project MNESYS (PE0000006) “A Multiscale integrated approach to the study of the nervous system in health and disease” (DN. 1553 11.10.2022). F.B. is partially supported by the MUR - PRIN project “Discrete random structures for Bayesian learning and prediction” no. 2022CLTYP4.
@@ A.1 Preliminary facts
We collect here some properties of multivariate normals, matrix normals and Wishart distributions to be used later. See Gupta and Nagar (2000) for details. In what follows $\operatorname{vec}(A)$ denotes the stacking of the columns of a matrix $A$ to form a vector.
@@
    Equivalence with the multivariate normal. $Z\sim\mathcal{MN}(0,\Sigma_{1},\Sigma_{2})$ if and only if $\operatorname{vec}(Z)\sim\mathcal{N}(0,\Sigma_{2}\otimes\Sigma_{1})$
@@
    Versions of this formula can be found in literature as Ingham-Siegel integrals (Ingham, 1933; Siegel, 1935, and  Fyodorov, 2002 for an extension to Hermitian matrices); in this generality, the statement can be derived by Proposition 8.3 and its proof in Eaton (2007). In particular the formula above holds for $\alpha>0$ and $S=\mathbb{1}_{D}$ and $C$ symmetric and $\geq 0$, showing that when $\mathbf{Z}_{j}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,C)$, then